# Walking the Tightrope of LLMs for Software Development: A Practitioners’ Perspective

### v1

@@ 3.2.1 Individual Level
TABLE IV: Summary of the Disadvantages of using LLMs at the Individual (Software Practitioner) Level.
Requirement Engineering &   Software Design Task: Brainstorming
[…]
Tasks: Diagram Generation
[…]
Software Development &   Software Quality Assurance Tasks: Information Retrieval
[…]
Tasks: Code Understanding
[…]
Task: Code Generation
[…]
Tasks: Test Generation; Test Data Generation; Code Translation; Pull Request Generation
[…]
Software Maintenance Tasks: Debugging
[…]
Tasks: Code Review
[…]
Increasing the Developer’s Effort. The majority of the participants (i.e., P1, P3-P12, P14-P17, P19, P21, P22) presented scenarios where adopting LLMs increase their effort, for example, due to LLMs’ response size limitation, e.g.: 9 “The biggest [limitation] would be the question of answer size, where it can’t answer a very large answer, and I’m going to have to ask several small questions in order to get my objective.” – P1 [Software Developer]. In order to avoid LLMs from getting a full picture, software developers may also need to break down prompts into subprompts, e.g.: 9 “I don’t typically give the whole idea that I want to achieve to the GPT. Rather, I kind of break down into different segments and then let them to generate the code for me for each segment. In the end, I would do my own construction [, integration,] of these segments.” – P17 [Data Analyst]. The unstable accuracy of LLMs also makes the developers spend energy in vain implementing wrong suggestions, e.g.: 9 “I would follow the debugging steps that they laid out for me. And then sometimes that works. Sometimes it doesn’t, as usual” – P3 [Software Engineer].
LLMs Dragging out Tasks. Many participants (i.e., P1, P3, P5-P8, P10-P11, P13-P17) mention how LLMs dragging out tasks. For instance, participants argue that there are situations where it is faster to do by themselves, e.g.: 9 “Sometimes you have [moments when you get doubtful about LLMs]: ‘Okay, I’m just struggling to get this done because [LLMs] it doesn’t get what I’m saying. So I’ll just do myself’” – P10 [Software Engineer].
Degrade Code Quality. Most of the participants (i.e., P1-P3, P5, P7-P11, P13-P22) discussed how LLMs negatively affect code quality. This can be caused by a weakening in their software development skills, e.g.: 9 “If you fully rely on it, the quality of your code base would drop by a lot” – P3 [Software Engineer]. They also highlight that LLMs may suggest solutions without taking into consideration the organisation’s approach, e.g.: 9 “Usually I make a lot of changes to the code as well, because there’s always something [that] the LLM will miss. The LLM won’t have the entire context of the task, or [understand] what’s the future plan [adopted by] the company” – P10 [Software Engineer]. LLMs can also provide wrong suggestions, e.g.: 9 “Sometimes ChatGPT - I don’t know if it’s [related to ChatGPT] design or something [else] - [recommend] libraries or packages [which] are not compatible. What ChatGPT proposed is not compatible with the environment I work in. I’ve encountered numerous times that I have to deal with this issue, and [I] ended up wasting a lot of time.” – P17 [Data Analyst].
Losing Learning Opportunities. Fourteen participants (i.e., P2-P8, P14-P15, P17, P19-P22) mentioned the negative impact on learning due to adopting LLMs. For instance, software developers might lose interesting discussions from online forums due to adopting LLMs, e.g.: 9 “from Stack Overflow, [there are] more comprehensive discussions there” – P8 [Software Engineer]. Relying on LLMs to reduce effort via automation may take away valuable learning opportunities.
Reducing Developer’s Code Mental Model. A few participants (i.e., P6, P10, P13-P14, P20-P21) mentioned the negative impact on the developer’s mental model. They argue that when developers overrely on LLMs by handing over the control to them, they lose in code understanding. This also affects software developers’ debugging capabilities, e.g.: 9 “For example, Copilot developed five class methods for me, I didn’t do it myself. If you debug it or show it to someone, you won’t know how the code was implemented.” – P6 [Data Engineer]. Consequently, they are pressured to continue relying on LLMs for debugging, e.g.: 9 “If you have zero knowledge of the code, you just wrote because ChatGPT wrote it, and you didn’t actually go through and read it, [doing code review], you then have to [rely on ChatGPT and to] ask ChatGPT to read the code that it wrote and then [get ChatGPT] tell it what problem happened” – P13 [Software Developer]. On the other hand, their code understanding improves naturally when coding by themselves, e.g.: 9 “When you code yourself, you already have that [mental model] by default, because you had to code.” – P10 [Software Engineer].
Negative Impact on Developers’ Personality. We found in the responses of thirteen participants (i.e., P1, P3-P5, P9, P11, P13, P15, P17, P19-P22) examples of how using LLMs impacts their personality. For instance, LLMs’ automation capabilities influence software developers in becoming lazier, e.g.: 9 “I have become lazy [to write] even if [it is just a] small one line I can also rely on the LLMs […] if I have to read a documentation, I will just ask LLMs to summarise so that I don’t have to read it completely” – P19 [Machine Learning Scientist], apathetic, e.g.: 9 “Even my capacity for analysis, it’s lost over time. I’ll spend more and more time just accepting, just like you review a PR (pull request). And you don’t really care” – P4 [Software Developer], and less confident in their own development skills, e.g.: 9 “It would take much longer to do something [that while not using AI, compared to] you [that] can do it with ChatGPT or Claude nowadays.” – P15 [Researcher].
Hindering Developers’ skills. Eleven participants (i.e., P2-P5, P7, P11-P15, P19-P20) mention how adopting LLMs affects software developers’ skills. This occurs because, when novice software developers overrely on LLMs to reduce effort by delegating tasks, they miss essential opportunities for the development of their technical and soft skills, e.g.: 9 “You’re so reliant on ChatGPT to think for you […] your skills don’t really increase much. You’re still that beginner” – P5 [Software Developer]. For more experienced developers, the process of atrophying their software development skills happens due to reliance on LLMs, e.g.: 9 “One of my colleagues said he deliberately turned off the [AI-based] completion because now he says ‘sometimes I just forget how it works. Sometimes I forget about the basics of my programming, because I’m used to this now’.” – P12 [Software Engineer], 9 “But since you’re missing that sort of mental exercise, you don’t really develop the coding muscles. And I think it kind of deteriorates your skills and your ability as a software developer.” – P5 [Software Developer]. Doing code review is not defended as not being enough to stop the skill deterioration, e.g.: 9 “‘Ah but [when using LLMs] you analyse the code’, I do. But I think that even my capacity for analysis, it’s lost over time. I’ll spend more and more time just accepting, just like [when] you review a PR (pull request) and you don’t really care. I think I’d do the same thing over time, gradually […] my review would get worse.” – P4.
Figure 4: Main Disadvantages of using LLMs at the Individual (Software Practitioner) Level.
Damaging Developers’ Reputation. A few participants (i.e., P6-P7, P9-P10, P19, P21-P22) mention how errors from code generated via LLMs can affect developers’ credibility - due to LLMs’ non-deterministic nature. For this reason, participants demonstrate a cautious approach towards completely delegating tasks to LLMs, e.g.: 9 “It [code with errors] impacts your credibility as a professional” – P9 [AI Engineer], 9 “When you push the code to the Git [repository], it is not the LLM’s name that will be in that log, it will be your name.” – P19 [Machine Learning Scientist].
Disrupting Software Development Flow. We found eleven participants (i.e., P4-P5, P7-P8, P12-P13, P15-P19) mentioned how interaction with LLMs can disturb their flow. For instance, interacting with LLMs via their online platform may cost a context switch to the code editor, e.g.: 9 “If I need to find a solution with it, I’m still [do] context switch because [ChatGPT] it’s not inside [the IDE] where I’m coding” – P16 [R&D Researcher]. The many suggestions offered by LLMs can also distract software developers, e.g.: 9 “Sometimes ChatGPT will give you an answer that’s completely different to what you were thinking. Unless you actually implement those [potential solutions] in real life, you don’t know what the result is going to be […] [basically] ChatGPT puts another option on top of your solution that you have in your mind.” – P7 [BI Analyst].
Communication Problems. We identified six participants (i.e., P1-P4, P6, P17) mentioning problems related to miscommunication with LLMs. This appears to happen due to AI hallucination or “misunderstanding” of user prompts, e.g.: 9 “Sometimes it will do extra things that I didn’t ask for.” – P5 [Software Developer]. But this can also occur due to the user facing difficulty in creating clear prompts, e.g.: 9 “There’s the communication problem. I can’t explain the whole robust part so that the LLM has the basis to be able to answer me” – P6 [Data Engineer].
3.2.2 Team Level
Losing Mentorship Opportunities. Some participants (i.e., P2-P3, P6-P7, P15-P14, P18-P19, P21-P22) describe how LLMs negatively influencing mentorship opportunities, taking the mentorship role from senior developers. This happens because novice developers would seek assistance first from LLMs, 9 “If I’m researching something that I don’t know a lot about, I usually will turn to technical documentation and Google […] first. While mentoring a junior in my field, they were very quick to go to LLMs first, and even criticise me, not for going to Google first” – P22 [Software Developer].
3.2.3 Organisation Level
Issues with Code License. A few participants (i.e., P2, P11, P16) mention license concerns at the organisation level, and how they shape their organisation’s approach towards LLMs, e.g.: 9 “Our company doesn’t allow the use of ChatGPT because of […] ownership [issues]. It’s not possible to get the ownership” – P18. Seeking to avoid potential code license issues, companies can decide to prohibit using LLMs in the workplace, e.g.: 9 “At my previous work […] their policy was just to avoid using ChatGPT or any other LLMs. Their concern was more to do with the ownership of the code that was being generated, and how it might violate the trademarks.” – P2 [Software Engineer]. Settling licenses may also cause a delay in the adoption of LLMs at the workplace, e.g.: 9 “We had to wait for the licenses [regarding ChatGPT] to come for that” – P16 [R&D Researcher].
Issues with Security & Privacy. We found security and privacy concerns in the responses of twelve participants (i.e., P2-P3, P5, P7, P9-P11, P13-P16, P18). There is a sceptical attitude on how LLMs approach these topics, e.g.: 9 “For security reasons, my current company isn’t allowing AI [to be] integrated into the program.” – P7 [Data Analyst]. The potential of LLMs to suggest code with potential security flaws also makes software developers approach LLMs with caution, e.g.: 9 “ When I started using it… I kind of realised that it [ChatGPT] tends to hallucinate a bit more. It made me worried about using the code directly in my code base. Because if I’m not careful and read every single line of code, I might introduce […] a security flaw in the code.” – P5 [Software Developer].
Cost of using LLMs. Seven participants (i.e., P2, P9-P10, P12, P16-P17, P19) mentioned the organisation cost of adopting LLMs. In this sense, the cost of tokens may limit high-volume interactions (i.e., vibe coding), e.g.: 9 “When we try to scale [that AI-generated code], we know that it becomes much faster and much more expensive quickly because we rely on OpenAI to do the calls, and we pay by token. You can’t actually do it because it’s going to cost this amount. – P16 [R&D Researcher]. This forces more planning before initiating this level of interaction, e.g.: 9 “We have to measure that in advance to make sure that you can spend that money [for] that type of stuff.” – P16 [R&D Researcher].
3.2.4 Society Level
Erosion of Social Trust. We found negative effects of LLMs on social trust in six participants (i.e., P2, P11, P14, P16, P19, P22). There is a concern of LLMs being misused, such as cheating during interview processes, e.g.: 9 “There’s also a lot of LLMs that can help you pass an interview, right? [Of course] it really depends on how [the interview] it’s monitored.” – P14 [Data Engineer], 9 “People would still need to consider the situation where it would add to misinformation and generate maybe a lot of fake news and fake videos […] [because of] that people would find it difficult to trust anything” – P2 [Software Engineer].
Job Market Crisis. Eight participants (P3, P7, P12, P16, P18, P20-P22) describe the impact of LLMs on the job market. Automation, not just LLMs, appears as the motivation for human replacement in simple tasks, e.g.: 9 “With the rise of AI and automation, a lot of things that humans can do can be replaced by using a robot.” – P3 [Software Engineer]. While there is still a demand for experts, non-specialised professionals might be replaced, e.g.: 9 “We’re still gonna need to rely on those experts for a lot of things. But then the issue is the ones that are not the experts, the in-between ones.” – P16 [R&D Researcher].
RQ2. How do LLMs hold software developers back? We found disadvantages at the individual, team, organisation, and society levels. At the individual level, LLMs may slow software development, increase developer’s effort, degrade code quality, reduce learning opportunities and developer’s code mental model, negatively impacts developers’ personality, hinder developers’ skills, damage developers’ reputation, and disrupt software development flow. Most of the participants mentioned LLMs slowing down software development and increasing developers’ effort.
3.3 RQ3. How do software developers achieve a balanced use of LLMs?
Exploring LLM Capabilities. This involves investigating, learning by experimentation, how LLMs can support software developers. In this process, developers can navigate through different LLMs until they find one that suits their needs, e.g.: 9 “I actually started with ChatGPT like everyone else. And then I found that ChatGPT [basically] give me too [much] general information […] then I tested [Microsoft] Copilot, and it gave me better results […] I think that, in some cases, it’s giving me not I wanted it and then I moved to Perplexity.AI, because Perplexity.AI has more current [information].” – P8 [Software Engineer]. Their perception towards LLMs also advanced during this process, e.g.: 9 “In the beginning, it seemed like they managed to make this semi-sentient chatbot thing. But as I started to understand the technology behind it, it became clear that it’s not very sentient. And it’s not really aware of what it’s doing. It’s more like a statistical or a machine learning model” – P5 [Software Developer].
Pragmatic Attitude towards LLMs. Participants describe slow adoption of LLMs. Some participants mention why the late adoption, e.g.: 9 “I just didn’t feel like I really needed [to use LLMs] […] I already have a good workflow established.” – P11 [Web Developer]. At the same time, software developers should drop LLMs in case they are not being helpful, e.g.: 9 “If you’re finding that it wastes your time, just stop using it and just write it yourself because [LLMs] it’s useful to save time. But if it’s not saving time, just stop using it.” – P13 [Software Developer].
Balancing Time-Saving & Learning Opportunities. While LLMs can provide time-saving for software development by reducing effort, that effort may be necessary to practice their skills. Software developers, especially novice developers, should seek a balance between time-saving and skill growth, e.g.: 9 “I’m going to continue seeing it [GenAI] as a tool that I can save me a lot of time. But at the same time, I can use it to help me learn as well.” – P14 [Data Engineer]. In that sense, it requires certain levels of self-control to avoid the temptation of overreliance on LLMs, e.g.: 9 “I use [Copilot] a good amount, but taking into account that I know I could use more, I just want to preserve the processing in my head […] If I don’t exercise my [brain for] programming logic, I’ll lose it [the skills].” – P4 [Software Developer].
Understand Suitable Use Cases for LLMs. Every participant presented suitable and unsuitable use cases for LLMs. LLMs are skilled in simple tasks and popular topics, e.g.: 9 “ChatGPT is very good at doing things that have already been done […] a lot of times.” – P5 [Software Developer]. Regarding unsuitable tasks, participants mention how LLMs are unhelpful for complex tasks, e.g.: 9 “If the tasks are fully related to code, design patterns, modularisation, or unit tests, yeah, [LLMs are] pretty helpful. But if it’s more related to the business aspect of the solution, LLMs are not that helpful.” – P9, or tasks that differ between companies, e.g.: 9 “When you are dealing with non-functional aspects of the solution, sometimes LLMs ask a lot about […] [for instance,] how to deploy this - because the deployment process [is] different among different companies.” – P9 [AI Engineer].
Combining Different LLMs for Different Tasks. Many participants (i.e., P2-P5, P8-P10, P12, P14, P16) defend that LLMs should be treated as a tool, which involves understanding their capabilities. Participants highlight how different LLMs (e.g., ChatGPT, Claude) are more efficient for different tasks, e.g.: 9 “I’ve been using ChatGPT for user-related tasks and requirement-related tasks but for code-related tasks, I’ve been using the Copilot.” – P12 [Software Engineer]. In this context, the traditional software engineering seems to transition into agentic software engineering, e.g.: 9 “Most applications will become agentic applications, using LLMs to take actions […] these new protocols which are emerging like the model context protocol (MCP) from Anthropic and A2A from Google will be able to improve this kind of communication between software and agent software” – P9 [AI Engineer]. This aligns with Hassan et al. [33] when they proposed the Agentic Software Engineering (SE 3.0).
LLMs for Code Improvement instead of Code Generation. Participants mention how LLMs aid them to improve their code, e.g.: 9 “usually what I do when I want to create the logic, I don’t just say [to LLMs]: ‘Generate this [without showing any example]’ [but] I actually start typing and creating the code. And I use more of the auto-complete [feature] of the LLMs, [than] the actual generation stuff.” – P10 [Software Engineer]. In this sense, the software developer maintains a balanced level of control over LLMs. In addition, participants mention avoiding using LLMs for code generation from scratch due to concerns involving the potential impact on their skills, e.g.: 9 “for code generation, I would be a bit more sceptical and maybe not use it very often. Because it will also affect your skills as well” – P2 [Software Developer].
Running LLMs Locally. Participants P9 and P22 both decided to run LLMs (i.e., Jan.AI, Llama, H2O Danube, and Mistral) locally via Ollama, Llama.cpp, and LM Studio as an approach to deal with privacy concerns in hosted LLM services. They believe maintaining privacy is worth the cost incurred, e.g.: 9 “I’m avoiding hosted services [and using LLMs locally] is largely around my concerns around privacy, and I do think there is a cost to privacy […] these companies that are hosting the services will train on your input […] It is not a private conversation.” – P22 [Software Developer].
RQ3. How do software developers achieve a balanced use of LLMs? We found recommendations focused on achieving a middle ground between benefits and disadvantages. They are based on exploration of LLM capabilities, understanding suitable use cases for LLMs, combining different LLMs for different tasks, a pragmatic attitude towards LLMs, and balancing time-saving with learning opportunities.
@@ 4.1 Implications for Practice
Gains do not come without cost. We presented in section 3.1 the benefits and 3.2 the disadvantages of using LLMs. Our findings align with the literature (e.g., [34, 35, 36, 37]) by showing the multiple facets regarding the impact of using LLMs. While LLMs can boost software development, software developers can be affected in the long term in case of overreliance. Those developers could find themselves in a scenario where they stop improving their skills. On the other hand, developers who decided to have self-control against overreliance will obtain more long-term career benefits due to their mature skills. When it comes to environmental cost, surprisingly, only two participants (P16, P22) demonstrated concerns about the energy consumption, overlooking water consumption [38]. Shi et al. [22] suggested that developers employ program-centric techniques, such as program pruning and grammar augmentation, to reduce the number of tokens.
Balanced control is the solution. We identified participants exercising various levels of control over Large Language Models (LLMs), ranging from high levels of control, such as self-restraint (e.g., turn off GitHub Copilot autocomplete), to low levels, exemplified by vibe coding [39]. While it may seem appealing to relinquish control over LLMs by generating extensive pieces of code, doing so can increase the risk of hidden bugs and errors. Although LLMs can assist with software development tasks, software developers must retain oversight. Examples of balanced control include using Test-Driven Development (TDD) in conjunction with LLMs and employing prompts across multiple LLMs. Mathews et al. [40] found that employing TDD with models like GPT-4 and Llama 3 leads to a higher success rate in solving programming challenges.
LLMs as Approach to Mitigate Interruptions. Several studies have delved into understanding how to improve developer experience (DevEx) [41, 42] by mitigating interruptions. Zuger et al. [43] argue that frequent interruptions for knowledge workers can lead to organizational costs due to incomplete tasks. To address this issue, they developed ”a physical traffic-light-like LED with an automatic interruptibility measure based on computer interaction data”. From their large-scale and long-term field study, they found that their solution reduced the interruptions of participants by 46%. In this sense, we identified that teams adopting LLMs may result in fewer interruptions.
LLMs Facilitating Developers Reaching the Flow. While analysing how LLMs influence software development flow in Section 3.1 and 3.2, we observe the great potential of LLMs for facilitating developers reach flow, aligned with the literature [44, 23]. Ritonumi et al. [45] describe flow state as: “The characteristics of being in flow include deep involvement in the activity (often described as intense and focused concentration), a merging of action and awareness, loss of reflective self-consciousness, a sense of control, an altered sense of time, and autotelic experience”. They also mention optimal challenge, challenge-skill balance, and immediate feedback as flow antecedents; and no distractions or interruptions and positive developer experience as facilitator factors to reach flow. LLMs can support provide a better developer experience (i.e., DevEx [46]), by reducing frictions for simple and repetitive tasks.
Different Effects on Novice and Experienced Developers. When adopting LLMs for code generation, developers move from actively coding to conducting code review [47]. While experienced developers face a loss of coding muscles, they can rely on muscle memory to go back to their best performance. On the other hand, novice developers did not build those muscles, and overreliance will make them stuck in their current level. On the other hand, software developers can improve their code review skills by reviewing LLMs’ suggestions, essential for senior roles.
LLMs as Champion for Prototyping. Participants described the high potential of LLMs for prototyping, which aligns with the literature (e.g., [48, 49]). Vibe coding paradigm appears to suit this context, focused on the development of a minimum viable product (MVP) to validate hypotheses about customer needs. They can quickly communicate the product proposals to the public [50].
With Great Power comes Great Responsibility. Participants demonstrated different perspectives involving code ownership when using LLMs. We observed less experienced participants showing insecurity in recognising their code ownership. However, code ownership emerges when developers act responsibly [51], ensuring the quality of the code. This aligns with the emerging regulations worldwide that emphasise the importance of human authorship. For instance, the U.S. copyright law333www.copyright.gov/ai requires “sufficient human contribution or control” over AI-generated code.
4.2 Directions for Future Work

### v2

@@ 3.2.1 Individual Level
TABLE IV: Summary of the Challenges of using LLMs at the Individual (Software Practitioner) Level.
Rationale: The visualisation allows developers to create a mental model of code generated using LLMs.
  • Challenge: LLM-induced code quality risks. Rationale: The visualisation allows developers to review the workflow used in the code generated via LLMs.
  • Challenge: Risk of losing learning opportunities. Rationale: LLMs can suggest more appropriate architectures, overlooked by developers. This allows developers to improve their understanding of software design.
    Adopt self-hosted LLM solutions. Participants P9 and P22 run LLMs (i.e., Jan.AI, Llama, H2O Danube, and Mistral) locally via Ollama, Llama.cpp, and LM Studio as an approach to deal with privacy concerns in hosted LLM services.
    • Challenge: Issues with security & privacy
    Rationale: Employing self-hosted solutions eliminates concerns regarding the organisation’s data privacy.
  • Challenge: Risk of losing learning opportunities. Rationale: Deploying self-hosted solutions allows developers to have a better understanding of LLMs’ limitations.
    Integrate solutions, not just LLMs, into the workflow. Participant P9 defends that Copilot is an example of an AI-integrated solution, which comprehends GitHub Actions and GitHub – popular continuous integration and continuous delivery and code versioning platforms. Participant P22 incorporated an AI-assisted code review by connecting Ollama into their GitLab server.
    • Challenge: Developer flow interruptions through LLM interaction
    Rationale: Employing integrated LLM solutions allows for more contextualised and accurate assistance, reducing the need for human intervention.
    Combine different smaller models. After experimenting with H2O.ai [^3] https://h2o.ai, Participant P9 suggests using Small Language Models as a solid alternative to LLMs, e.g.: “In some parts of the solution, it’s possible to use smaller models very efficiently and diminish the cost of the overall solution” – P9 [AI Engineer].
    • Challenge: Cost of using LLMs
    Rationale: Employing smaller language models rather than large ones can help reduce computational, memory, and energy requirements [2].
    Prohibit LLMs from merging code into the main branch. Human in the loop is still required for accountability, e.g.: “If I’m putting up a pull request for my company, I need to understand that code. I need to have the ownership of it. Because if something fails, I won’t be blaming Cursor, or anyone else, this [error] will be on me to fix. Any code that’s generated by the LLMs, I only put this code in production [environment] if I actually understand that [code]” – P10 [Software Engineer].
    • Challenge: LLM-induced code quality risks
    Rationale: Implementing mandatory human review for main-branch commits can safeguard against errors in LLM-generated code [53].
  • Challenge: Reputational vulnerability through LLM delegation. Rationale: The credibility of developers and their organisations is preserved by safeguarding against errors from code generated via LLMs.
    Follow the established coding conventions. Participant P11 described how GitHub Copilot can get confused according to the used variables, e.g.: “When you have some kind of misleading or not super standard variable names, Copilot gets confused” – P11 [Web Developer].
    • Challenge: Frictions in human-LLM communication
    Rationale: By adhering to code conventions consistent with those found in the data used to train LLMs, developers can reduce the risk of miscommunication with these models.
  • Challenge: LLM-induced code quality risks. Rationale: By adhering to code convention best practices, code generated by LLMs is more likely to achieve satisfactory quality.
  • Challenge: Reputational vulnerability through LLM delegation. Rationale: By mitigating errors from code generated using LLMs, the credibility of developers and their organisations is preserved.
  • Challenge: Erosion of developer skills and agency. Rationale: By following code conventions, developers.
    Introduce LLM4SE usage policies. Four participants (18.1%) indicated in the pre-interview questionnaire that their organisations either do not have an AI policy or are not aware of one. Having a clear policy addressing the use of software development forms an essential foundation for effective IT governance.
    • Challenge: Issues with security & privacy
    Rationale: LLMs4SE policies that indicate the supported LLMs tools and how to share code with those tools may mitigate concerns regarding privacy.
  • Challenge: Issues with code license. Rationale: Defining how to handle code licensing in the LLM4SE policy can guide developers and help prevent potential legal issues in the future.
  • Challenge: Erosion of social trust. Rationale: Clarifying how LLMs manage developers’ prompts may help ease their concerns.
    Continuously upskilling in the latest SE practices and tools. Participants P1, P4, P10, P14, and P20 defend the importance of upskilling to not get behind in the market, e.g.: “I will always try to upskill myself in a way that I won’t get replaced by LLMs because the implementations are not perfect” – P20 [Software Engineer].
    • Challenge: Job market crisis
    Rationale: Keeping up-to-date with the latest SE best practices and tools, such as spec-driven development, may help developers to maintain relevance in this AI era.
@@ 4.1 Implications for Practice
    We analyse the implications of adopting LLMs in SE by exploring both the benefits (RQ1) and challenges (RQ2) using Marshall McLuhan’s Tetrad of Media Effects [41]. This framework enables us to examine how new technologies can augment human capabilities through different aspects. The Tetrad consists of four fundamental questions: What does the technology enhance or intensify? What does it make obsolete or displaced? What does it retrieve or recover from the past? And what does it reverse into when taken to extremes?
    Enhance. This explores how adopting LLMs enhances or intensifies practices and processes in software development.
    • Cognitive debt: In section LABEL:subsubsec:incompleteModel, we highlight how adopting LLMs can deprive developers of a complete mental model of the code. According to Storey [52], this cognitive debt may silently build up over time and lead to developers losing control over their software. In our study, participants keep control over LLMs by (i) relying on them for code improvement rather than code generation and (ii) employing Test-Driven Development (TDD). Mathews et al. [39] found that employing TDD with models like GPT-4 and Llama 3 leads to a higher success rate in solving programming challenges.
    • Company onboarding: In section 3.1.2, we present the potential of LLMs being incorporated into the onboarding process. However, in the ongoing case study by Azanza et al. [3] at an IT consulting organisation, they found that using third-party LLMs can pose unacceptable data risks for companies. They explored open-source LLMs for onboarding. Additionally, organisations may seek agreement with LLM vendors for data protection.
    Obsolete. This aspect explores the obsolescence or displacement caused by adopting LLMs for software development.
    • Repetitive tasks: In section 3.1.1, we mention how LLMs reduce developers’ effort for simple tasks, such as boilerplate code generation and information retrieval. However, it is still essential to keep humans in the loop, at least as reviewers, since LLMs are based on stochastic models.
    Retrieve. This examines how far adopting LLMs makes it possible to retrieve or recover concepts, practices, or processes adopted in the past in SE.
    • Exploratory learning: In section 3.1.1, we highlight how participants use LLMs as a learning tool, augmenting the traditional problem-solving process. Using the chat mode, similar to Hallway talks, developers can ask questions, simulating hypothetical scenarios, getting an independent perspective. We suggest that other developers be added to the loop for sharing knowledge purposes.
    Reverse. This explains how adopting LLMs for software development, when pushed to its extreme, undergoes a reversal or transformation.
    • Skill atrophy: In section LABEL:subsubsec:skillAtrophy, we highlighted that adopting LLMs may hinder developers’ skills. While LLMs can provide time-saving for software development by reducing effort, that effort is also necessary to practice their skills. In our study, participants restrict themselves from overusing AI (e.g., turning off AI-based suggestions once in a while) and force themselves to use their coding muscles.
    • Environmental crisis: When it comes to environmental cost, surprisingly, only two participants (P16, P22) demonstrated concerns about energy consumption, overlooking water consumption [25]. Shi et al. [50] suggested that developers employ program-centric techniques, such as program pruning and grammar augmentation, to reduce the number of tokens.
    4.2 Implication for Research